# REF-VLM: Triplet-Based Referring Paradigm for Unified Visual Decoding

### v1

@@ E.1 Task Definition
Table 19: Comparison of Referring Expression Comprehension (REC) performance. REC is evaluated by IOU@0.5 Accuracy. VGM represents vision generalist model and MLLM represents multimodal large language model.
@@
Table 20: Comparison of Referring Expression Segmentation (RES) performance. The “Decoder” column refers to the visual decoder utilized by the MLLM for performing RES tasks. indicates that the visual decoder is custom-designed and trained from scratch. The performance of RES is evaluated using cumulative IoU (cIoU) as proposed by [32].
@@
[…]
@@
[…]
Appendix F More Results
F.1 More Experimental Results
Referring Expression
For the Referring Expression Segmentation (RES) task, we evaluate REF-VLM on the RefCOCO, RefCOCO+, and RefCOCOg test and validation datasets by calculating the cumulative IOU (cIOU) as proposed by [32]. REF-VLM with meta plugins, trained from scratch, achieves results in both zero-shot and fine-tuned settings that are comparable to recent methods like LISA [28], which utilized pretrained backbones such as SAM (see Table 20). The results show that the performance of our REF-VLM trained from scratch is slightly lower than that of PixelLM [46], primarily due to PixelLM’s use of the CLIP-ViT-H visual encoder, which has significantly more parameters than ours (see Table 13). Additionally, to demonstrate that REF-VLM is not only capable of using custom-designed components but also can extend to current VGMs, we employed SAM as an external plugin for our mask decoder. By loading the pretrained weights from SAM and fine-tuning it on our VT-Instruct datasets, we found that REF-VLM with the external plugin outperforms current MLLMs such as VisionLLMv2 [56], GLaMM [45], and demonstrates comparable performance to the Generalist Model such as UNINEXT-H [63].
For the Referring Expression Comprehension (REC) task, we compare our REF-VLM with current MLLMs capable of generating referring boxes based on specific prompts in both zero-shot and fine-tuned settings. The metric used for REC evaluation is IoU@0.5. As shown in Table 19, REF-VLM demonstrates superior performance in the REC task compared to other MLLMs.
Interactive Grounding
For this task, we evaluate using the prompt, “Please generate a mask based on the region <region> in the image <image>.” where <region> is replaced with visual prompts such as points, scribbles, boxes, or masks. The results presented in Table 18 show that our REF-VLM with meta plugins outperforms both SAM [26] and SEEM-B [80] across point, scribble, box, and mask settings, achieving performance comparable to PSALM, which utilizes pretrained Swin-T and Mask2Former weights in these configurations. Furthermore, our REF-VLM with external plugins achieves superior performance compared to VisionLLMv2 [56] and PSALM [73].
Table 21: Comparison of different configurations of VPT encoders on RefCOCOg validation dataset.
Method CIDEr Meteor Params
Osprey [68] 78.5 12.0 6.55M
Use Mask Token 77.9 12.0 1024
No Mask Token 78.8 12.1 0
Table 22: Comparison of different configurations of Matching Strategy on COCOREM Test dataset.
Method AP50 AP75 AP[50:95] AR[50:95]
No-Matcher 12.4 1.4 4.0 16.4
Use Matcher 12.7 2.6 4.5 19.3
F.2 More Ablation Results
Different Configuration of VPT Encoding Strategy. Our parameter-free VPT encoder configuration outperforms alternatives as shown in Table Table 21. Without mask tokens, our approach achieves the best CIDEr (78.8) and Meteor (12.1) scores with zero additional parameters, using block-wise dot products between masks and visual features with position encoding and shared projectors. Conversely, Osprey [68] uses unblocked dot products (N=1) and separate projections, requiring 6.55M parameters while yielding inferior results. Our efficient design delivers both superior performance and exceptional cross-architecture adaptability.
Ablation of Group Hugrain Matcher
We conduct experiments with Group Hungarian Matching on the more challenging task of freeform object detection. Both experiments are trained for 5 epochs on the COCO-REM dataset, with comparative results shown in Table 1. The Group Matcher demonstrates superior performance across AP and AR metrics, particularly showing significant improvement in AP75, indicating its effectiveness in enhancing box regression accuracy.
Appendix G More Visualization Results
Prompt Provide a simple description of this photograph. I would like to hear a more thorough description. [Uncaptioned image] VD-CoT      <Task>      Unit decode (False).      </Task> Answer with Triplets A group of people, including children, are gathered on a brick wall, some holding umbrellas and backpacks while others wear dresses, sunhats, and shirts. A man wearing a hat and a pink-shirted person are also visible. In the background, a large brick building can be seen, and the scene outside provides a view.
Table 23: An Example of VD-CoT Applied to the Visual Caption Task.In this task, only textual responses are required, and no operations on image information are needed. Therefore, in the generation of VD-CoT, Unit decode is set to False, and no visual-related special tokens are generated. Similarly, in the result of Answer with Triplets, there are no decoded triplets related to visual content.
Prompt Please identify each object in the image. [Uncaptioned image] VD-CoT      <Task>      Unit decode (True). Class name, target unit and number:       - Name: car Unit: box Num: 10       - Name: skyscraper Unit: box Num: 1       - Name: barricade Unit: box Num: 1       - Name: city Unit: box Num: 1       - Name: street Unit: box Num: 1      </Task> Answer with Triplets <Phrase>car</Phrase> (<Unit>box</Unit>[0]<REF>[1]<REF>[2]<REF>[3]<REF> [4]<REF> [5]<REF>[6]<REF>[7]<REF>[8]<REF>[9]<REF>), <Phrase>skyscraper</Phrase> (<Unit>box</Unit>[0]<REF>), <Phrase>barricade</Phrase> (<Unit>box</Unit>[0]<REF>), <Phrase>city</Phrase> (<Unit>box</Unit>[0]<REF>), <Phrase>street</Phrase> (<Unit>box</Unit>[0]<REF>).
Table 24: An Example of VD-CoT Applied to the Object Detection Task. VD-CoT analyzes the visual content and generates decoding triplets to identify objects in the image. The special tokens and references have been included for clarity.
Prompt Delineate areas of the image into semantic segments. [Uncaptioned image] VD-CoT      <Task>      Unit decode (True). Class name, target unit and number:       - Name: couch Unit: mask Num: 1       - Name: computer keyboard Unit: mask Num: 1       - Name: mask Unit: mask Num: 2       - Name: person Unit: mask Num: 2       - Name: sofa Unit: mask Num: 1       - Name: monitor (computer equipment) computer monitor Unit: mask Num: 1      </Task> Answer with Triplets <Phrase>couch</Phrase> (<Unit>mask</Unit>[0]<REF>), <Phrase>computer keyboard</Phrase> (<Unit>mask</Unit>[0]<REF>), <Phrase>mask</Phrase> (<Unit>mask</Unit>[0]<REF>[1]<REF>), <Phrase>person</Phrase> (<Unit>mask</Unit>[0]<REF>[1]<REF>), <Phrase>sofa</Phrase> (<Unit>mask</Unit>[0]<REF>), <Phrase>monitor (computer equipment) computer monitor</Phrase>       (<Unit>mask</Unit>[0]<REF>).
Table 25: An Example of VD-CoT Applied to the Open-Vocabulary Segmentation Task. VD-CoT analyzes the visual content and generates decoding triplets for semantic segmentation of the image. Special tokens and references are provided for clarity.
Prompt Please pinpoint the egg yolk in the image and provide the dimensions of its bounding box. [Uncaptioned image] VD-CoT      <Task>      Unit decode (True). Class name, target unit and number:       - Name: target Unit: box Num: 1      </Task> Answer with Triplets <Phrase>target</Phrase> (<Unit>box</Unit>[0]<REF>).
Table 26: An Example of VD-CoT Applied to the Referring Expression Comprehension Task. VD-CoT analyzes the visual content and generates decoding triplets to identify the target (egg yolk) and its bounding box dimensions in the image.
Prompt How do you create a segmentation mask for a white boat in the water in the image? [Uncaptioned image] VD-CoT      <Task>      Unit decode (True). Class name, target unit and number:       - Name: target Unit: mask Num: 1      </Task> Answer with Triplets <Phrase>target</Phrase> (<Unit>mask</Unit>[0]<REF>).
Table 27: An Example of VD-CoT Applied to the Referring Expression Segmentation Task. VD-CoT analyzes the visual content and generates decoding triplets to identify the target (white boat) and creates its segmentation mask in the image.
Prompt What defining features can you describe for the area [VPT] in the image? [Uncaptioned image] VD-CoT      <Task>      Unit decode (False).      </Task> Answer with Triplets The cabinet is closed.
Table 28: An Example of VD-CoT Applied to the Referring Expression Generation Task. VD-CoT analyzes the visual content and generates descriptive features for the given area [VPT] in the image.
Prompt Segment the region based on the region [VPT] in the image, and illustrate the resulting mask. [Uncaptioned image] [Uncaptioned image] VD-CoT      <Task>      Unit decode (True). Class name, target unit and number:       - Name: boat Unit: mask Num: 1      </Task> Answer with Triplets <Phrase>boat</Phrase> (<Unit>mask</Unit>[0]<REF>).
Table 29: An Example of VD-CoT Applied to the Interactive Grounding Task. VD-CoT analyzes the visual content, segments the region [VPT], and generates a mask for the identified object (boat) in the image.
Figure 9: The Visual Understanding Results of REF-VLM.
Figure 10: The Detection Results of REF-VLM. The text color in the model’s responses corresponds to the bounding box colors of the detected objects in the images. For example, in the top-right image, the model detects two categories: “person” and “dress.” The “person” category contains two instances, represented by [0]<REF> and [1]<REF>, while the “dress” category contains one instance, represented by [0]<REF>.
Figure 11: The Segmentation Results of REF-VLM. The figure illustrates the segmentation and GCG segmentation outputs generated by REF-VLM. The text corresponds to the mask colors of the segmented objects in the images. For example, in the top-left image, the model segments three categories: “stool,” “person,” and “grassy field.” The “person” category contains two instances, represented by [0]<REF> and [1]<REF>, while the “stool” category contains one instance, represented by [0]<REF>. The background, labeled as “grassy field,” is also represented by [0]<REF>.
Figure 12: The Grounding Detection Results of REF-VLM. We use the textual prompts and visual box prompts for grounding detection tasks. In the first row (left), a textual prompt instructs the model to locate the object “bear” mentioned in the sentence. The second row displays the result, where the model successfully detects the location of the bear using a bounding box. In the first row (right), a bounding box around a bird is provided as a visual prompt. The textual query uses the special token <area> to refer the boxed region in the image. The second row (right) shows the model’s output, correctly identifying the object in the boxed region as a “bird.”
Figure 13: The Grounding Segmentation Results of REF-VLM. We use the textual prompts and visual box prompts for grounding segmentation tasks. In the first row (left), a textual prompt instructs the model to segment the middle bird in the center of the image. The second row (left) shows the segmentation result for the middle bird produced by the model based on the textual prompt. In the first row (right), bounding boxes around two birds are provided as visual prompts, and the textual prompt includes a special symbol <area> to indicate the region for segmentation. The second row (right) displays the segmentation results for the left bird and right bird based on two types of prompts.

### v2

@@ V-A Quantitative Results
TABLE VII: Comparison of Referring Expression Comprehension (REC) performance. REC is evaluated by IOU@0.5 Accuracy. VGM represents vision generalist model and MLLM represents multimodal large language model.
@@
TABLE VIII: Comparison of interactive grounding performance on segmentation task. The task is evaluated on the COCO-Interactive [69] validation dataset. The evaluation metrics for interactive grounding are mIoU and cIoU.
Model Decoder Scratch Type Point Scribble Box Mask
mIoU cIoU mIoU cIoU mIoU cIoU mIoU cIoU
SAM-B [26] - VGM 48.7 33.6 - - 73.7 68.7 - -
SAM-L [26] - 51.8 37.7 - - 76.6 71.6 - -
SEEM-B [76] - 47.8 57.8 43.0 44.0 44.9 42.1 48.4 65.0
PSALM [69] Mask2Former
MLLM 64.3 74.0 66.9 80.0 67.3 80.9 67.6 82.4
VisionLLMv2 [54] GroundingDINO
65.4 70.9 66.8 77.2 74.2 83.2 67.9 83.8
REF-VLM (meta) Mask Decoder 62.8 70.4 59.8 60.2 71.2 73.7 66.3 77.5
REF-VLM (external) SAM
65.6 75.2 68.3 79.4 74.9 84.6 68.2 83.7
TABLE IX: Evaluation on zero-shot open-vocabulary tasks. Results on ADE20K (instance segmentation) and COCO2017 (object detection). As no existing MLLM supports zero-shot detection without class prompts, only the proposed model is evaluated.
Model Type Scratch
Mask
Decoder
Box
Decoder
ADE20k COCO
MaskCLIP [18] SEG MaskCLIP - 6.0 -
ODISE [56] SEG Diffusion UNet - 14.4 -
SAN [58] SEG CLIP+SAN - 10.6 -
PSALM [69] SEG
Mask2Former - 9.0 -
PSALM+LVIS [69] SEG
Mask2Former - 13.9 -
REF-VLM ($\text{mAP}_{S}$) SEG/DET MaskFormer DETR 16.7 26.7
TABLE X: Comparison of Referring Expression Segmentation (RES) performance. The “Decoder” column refers to the visual decoder utilized by the MLLM for performing RES tasks. indicates that the visual decoder is custom-designed and trained from scratch. The performance of RES is evaluated using cumulative IoU (cIoU) as proposed by [32].
@@
PixelLLM [57] SAM
78.5 74.4 76.9 72.1 64.5 69.2 72.4 70.7
AnyRef SAM
79.9 74.2 76.9 73.5 61.8 70.3 70.7 70.0
NExT-Chat [66] SAM
78.9 69.5 74.7 71.9 56.7 65.1 67.0 67.0
VITRON [19] SEEM
78.7 71.6 74.4 72.1 57.8 66.3 67.3 67.2
GroundHOG [68] Mask2Former
79.9 75.7 78.5 75.0 64.9 70.5 74.6 74.1
GLaMM [43] SAM
83.2 76.9 79.5 78.7 64.6 72.6 74.9 74.2
VisionLLMv2 [54] GroundingDINO
82.3 77.0 79.2 75.8 61.8 68.9 74.8 73.3
@@
REF-VLM (external) SAM
82.9 76.8 81.2 77.6 63.4 73.1 75.0 74.6
TABLE XI: Comparison across visual encoder and ablation study on group matchers. [0,1,2,3]: all CLIP-ConvNeXt layers; -2: second-to-last CLIP-ViT layer; [0,1,2,4]: first three CLIP-ConvNeXt layers plus CLIP-ViT output. ✔indicates the use of the Group Hungarian Matcher during decoder training.
Visual Encoder Size
Group
Matcher
Feature
Dimension
cIoU
ConvNeXt-L 320 [0,1,2,3] 41.94
ConvNeXt-L 336 [0,1,2,3] 46.08
CLIP-ViT 336 -2 60.44
ConvNeXt-L + CLIP-ViT 320 [0,1,2,4] 60.02
ConvNeXt-L + CLIP-ViT 512
[0,1,2,4] 61.83
ConvNeXt-L + CLIP-ViT 512 [0,1,2,4] 62.49
Freeform Open-Vocabulary Identification. Compared to existing MLLMs, which generally require category prompts to be added in open-vocabulary tasks [54, 69], enabling the model to understand potential categories in the image before performing open-set detection or segmentation, our REF adopts a more flexible approach. With REF-VLM, there is no need to predefine categories, we simply input the prompt like “Please detect bounding boxes (segment objects) in the image<image>”. With this prompt, REF-VLM can autonomously identify potential categories within the image through its LLM, handing off decoding tasks to the subordinate visual decoder. We refer to this more adaptable task as Freeform Open-Vocabulary Identification. Unlike other MLLMs [54, 69], REF-VLM allows the LLM to directly output categories instead of having the downstream decoder output category names and confidence scores. Accordingly, instead of the calculating mAP as the metric for Open-Vocabulary Identification tasks, we propose a new metric called mAP Similarity ($\text{mAP}_{S}$) to evaluate our REF-VLM performance. For traditional open-vocabulary models, they typically predict classes with a logit score by their classification head. However, instead of applying a classification head for each task, our REF-VLM leverages a large language model (LLM) to predict classes without generating any class logits. We therefore compute the similarity score between REF-VLM’s class predictions and all ground truth class names. We then assign the class label based on the highest similarity, using this similarity score in place of the traditional confidence score.
For the implementation of $\text{mAP}_{S}$, we denote the phrases predicted by the LLM as $\mathcal{P}=\{p_{1},p_{2},\dots,p_{k}\}$, where $k$ is the number of predicted phrases. The ground-truth class set is defined as $\mathcal{C}=\{c_{1},c_{2},\dots,c_{n}\}$, where $n$ denotes the total number of classes in the dataset. We employ the CLIP-Large-14-336 model to encode both predicted phrases and ground-truth class names into a shared embedding space, yielding text embeddings $\mathbf{e}_{p_{i}}=f_{\mathrm{CLIP}}(p_{i})$ and $\mathbf{e}_{c_{j}}=f_{\mathrm{CLIP}}(c_{j})$, respectively. For each predicted phrase $p_{i}$, we compute the cosine similarity between its embedding and all class embeddings as $s_{i,j}=\frac{\mathbf{e}_{p_{i}}^{\top}\mathbf{e}_{c_{j}}}{\lVert\mathbf{e}_{p_{i}}\rVert_{2}\lVert\mathbf{e}_{c_{j}}\rVert_{2}}$. The predicted class for $p_{i}$ is then determined by $j^{*}=\arg\max_{j}\;s_{i,j}$, and the corresponding maximum similarity score $s_{i,j^{*}}$ is used as the logit score for evaluation.
We evaluate REF-VLM on the ADE20k test dataset for zero-shot freeform open-vocabulary segmentation task and COCO2017 validation dataset for object detection. The results for both tasks are presented in Table IX. Notably, REF-VLM achieves strong performance without any specialized design, outperforming other MLLMs (e.g., PSALM [69]) and specialist models (e.g., SAN [58]). Additionally, REF-VLM also demonstrates the capability to perform open-vocabulary object detection.
Significance Analysis of TRP. We selected a general segmentation task with the prompt “please segment objects in the image <image>.” The model outputs two classification results: “a bride” and “a giant illuminated heart”, using our designed VD-CoT format. We used attention visualization techniques to generate attention significance maps for the REF token with respect to each output token of the MLLM. The attention values from each layer were averaged and visualized, as shown in the figure above. The Fig. 6 demonstrates that each <REF>token exhibits high attention response values to the preceding <Phrase>, <Unit>, and output numbers.
Figure 6: Attention Significance of <REF>Tokens for MLLM Output Tokens. Under the TRP paradigm, the generated prior information guides the formation of decoding features.
V-B Ablation Study
Choose of Group Matchers. To validate the effectiveness of our Group Hungarian Matcher, we perform an ablation study on its usage in the mask decoder for the RES task, using the RefCOCOg test dataset and cIoU as the evaluation metric. As shown in Table XI, applying the Group Hungarian Matcher for loss computation yields a significantly better performance compared to configurations without it, demonstrating its substantial impact on improving the overall accuracy.
Configuration of Visual Encoders. To investigate the effect of different configurations of CLIP vision encoders, including CLIP-ConvNeXt and CLIP-ViT, along with variations in image size and feature selection layers, we conduct experiments on the RES task using the RefCOCOg test dataset. As shown in Table XI, REF-VLM achieves the highest performance when concatenating the CLIP-ConvNeXt and CLIP-ViT encoders with the setting of image size as 512×512.
VI Conclusion
In conclusion, we propose REF-VLM, an extensible open-ended visual multi-task learning framework, together with TRP, a unified referring paradigm, and VT-Instruct, a large-scale multimodal instruction-tuning dataset. Extensive experiments demonstrate the effectiveness of REF-VLM across diverse visual tasks and settings. Nonetheless, current task coverage remains limited, and performance in multi-turn dialogues as well as data utilization require further investigation. Future work will focus on expanding supported tasks, improving multi-turn robustness, and optimizing training data usage efficiency.